of meters. 30. Given these two vectors A (3.00)i + (7.00)j - (4. 00) k and B (2.00)i - (4.50)j + (3. 00) k: a. Calculate A. B Calculate A >< B b. Calculate the angle between A and B Determine by calculation if the vector c. is perpendicular to B. d. Calculate a unit vector that points in the direction of the vector-A
Answer #1

a)

A . B = (3 × 2)-(7 × 4.5) _ (4 × 12) = 6-31.5+ 12 -37.5

2 -4.5 3

b) Angle between A and B is

-37.5 COS-1 AI B cos-1 (-0.756)

Hence

θ-139. 110

c) If dot product of B and C is zero, then they are perpendicular

B. С= (2x 7.8) _ (4.5 × 4.6) _ (3x 1.7) =-10.2

Therefore C is not perpendicular B

d) Unit vector in the direction of -A is

ー-03492-0.814j 0.465k 1시 74